AdiGirl:

Joe walked four over nine of a mile in four over five of an hour. What is his unit rate in miles per hour? four ninths over four fifths = five over nine mile per hour four fifths over four ninths = one and four fifths miles per hour four ninths over four fifths = one and four fifths miles per hour four fifths over four ninths = five over nine mile per hour

AdiGirl:

@Shadow

AdiGirl:

can you help me?

Shadow:

So first you want to convert to the same denominator. Right now you have 4/9mile in 4/5hour. \[\frac{ 4 }{ 9 } \times \frac{ 5 }{ 5 } = \frac{ 20 }{ 45 }\] \[\frac{ 4 }{ 5 } \times \frac{ 9 }{ 9 } = \frac{ 36 }{ 45 }\] 20/45mile in 36/45hour \[(\frac{ 20 }{ 45 } \div \frac{ 36 }{ 45 }) = \frac{ 20 }{ 45 } \times \frac{ 45 }{ 36 } = \frac{ 20 }{ 36 } = \frac{ 5 }{ 9 }mph \]

AdiGirl:

ok

Shadow:

Just analyze what I wrote and you should be able to find the answer.

AdiGirl:

im confused

AdiGirl:

do I have to multiply 20/45 and 45/36 orr divide it

AdiGirl:

um @Shadow

bm717:

When you divide two fractions, you must multiply the second fraction by the reciprocal. Sorry he's a lazy person.

bm717:

\[\frac{ a }{ b }\div \frac{ c }{ d }~ is~the~same~thing~as~\frac{ a }{ b } \times \frac{ d }{ c }\]

AdiGirl:

ohhhhhhhhh got it

bm717:

You must flip the second fraction.

bm717:

Then you multiply across and he got the answer \[\frac{ 5 }{ 9 }\]

AdiGirl:

so I just gotta find the same answer but from the answer choices?